OpenStudy (anonymous):

Which of the following ordered pairs lies on the graph of h(x) = -2x2? (1, -4) (1, -2) (1, 4)

OpenStudy (anonymous):

try x value in the equation and see if you get h(x)

OpenStudy (anonymous):

(1,-4) h(x)=-2x^2 h(1)=-2(1)^2 h(1)=-2 this value does not much the pair (-4) so this pair is wrong

OpenStudy (anonymous):

so its not b so it must be a right?

OpenStudy (anonymous):

is not (1,-4)

OpenStudy (anonymous):

ohh :O

OpenStudy (anonymous):

try (1,-2)

OpenStudy (anonymous):

kk :3

OpenStudy (anonymous):

nvm i turned in the assignment

OpenStudy (anonymous):

h(x)=-2x^2 h(1)=-2(1)^2=-2 so the pair is (1,-2)